Suppose a U.S. investor wants to invest in a British firm currently selling for ₤50 per share. The investor has $7,000 to invest, and the current exchange rate is $1.40/₤.

After 1 year, the exchange rate is unchanged and the share price is ₤55. What is the dollar-denominated return?

A. 14% B. 10% C. 9.3% D. 7.1%
Answer #1

Exchange Rate = $1.40/₤

Value of one share now = ₤50 = $50*1.40 = $70

Value of share after one year = ₤55 = $55*1.40 = $77

Return = (Value of share after one year - Value of the share now)/Value of the share now * 100%

= (77 - 70)/70 *100 = 10%